For normal distribution with mean zero and variance σ2 show that:


E(|x|)= (√2/x)σ



1
Expert's answer
2021-06-07T18:47:58-0400

The probability density function (PDF) for a normal "X\sim (\mu, \sigma^2 )" is:


"f_X(x)=\dfrac{1}{\sigma \sqrt{2\pi }}e^{-\frac{1}{2}(\frac{x-\mu}{\sigma})^2}"


For normal distribution with mean zero and variance σ2

"f_X(x)=\dfrac{1}{\sigma \sqrt{2\pi }}e^{-\frac{1}{2}(\frac{x}{\sigma})^2}"

"E(|X|)=\displaystyle\int_{-\infin}^{\infin}|x|f_X(x)dx"




"=2\displaystyle\int_{0}^{\infin}x\dfrac{1}{\sigma \sqrt{2\pi }}e^{-\frac{1}{2}(\frac{x}{\sigma})^2}dx"

Let "z=\dfrac{x}{\sigma}"

Then "dx=\sigma dz"


"E(|X|)=\displaystyle\int_{-\infin}^{\infin}|x|f_X(x)dx"

"=2\displaystyle\int_{0}^{\infin}x\dfrac{1}{\sigma \sqrt{2\pi }}e^{-\frac{1}{2}(\frac{x}{\sigma})^2}dx"

"=\displaystyle\int_{0}^{\infin}\dfrac{2\sigma}{ \sqrt{2\pi }}ze^{-\frac{1}{2}z^2}dz"

"=-\dfrac{2\sigma}{ \sqrt{2\pi }}\big[e^{-\frac{1}{2}z^2}\big]\begin{matrix}
 \infin \\
 0
\end{matrix}=\sqrt{\dfrac{2}{\pi}}\cdot\sigma"

"E(|X|)=\sqrt{\dfrac{2}{\pi}}\cdot\sigma"